OpenStudy (anonymous):

The volume of a sphere with radius r is given by the formula v=4/3 πr3 Find the volume of a sphere with radius 4 meters. Use 3.14 for the value of π and round your answer to the nearest tenth.

OpenStudy (anonymous):

\[\Large V=\frac{4}{3}(\pi )r^3\] use r=4 and pi=3.14 \[\Large V=\frac{4}{3}(3.14)(4)^3\] use your calculator and let me know answer .

OpenStudy (anonymous):

what is pi

OpenStudy (anonymous):

\[\pi=3.14\]

OpenStudy (anonymous):

37.68

OpenStudy (anonymous):

nopes ! check it again.

OpenStudy (anonymous):

3.14*4=12.56 4*3=12 12+12.56=24.56

OpenStudy (anonymous):

so v=24.56

OpenStudy (anonymous):

ok fallow the steps first take cure of 4 =4*4*4=64 then multiply with 3.14 64*3.14=200.96 now multiply with 4 200.96*4=803.84 now divide with 3 \[\Large \frac{803.84}{3}=267.94\]

OpenStudy (anonymous):

cube*

OpenStudy (anonymous):

ooo oki thougt 4^3 means 4*3

OpenStudy (anonymous):

so its ok now ?

OpenStudy (anonymous):

sorry so would i put i as v=267.94

OpenStudy (anonymous):

yes of course

OpenStudy (anonymous):

ok thankyou so much

OpenStudy (anonymous):

yw:)